# Response of superfluid fermions at finite temperature

### v3

@@ II.1.2 Finite temperature
^${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=-\langle{\cal T}{\hat{\Psi}}_{\mu\mu^{\prime}}(\tau){\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}(\tau^{\prime})\rangle$ (80)
@@
The operator ${\cal H}$ is the redefined Hamiltonian ${\cal H}=H-\lambda N$, where $\lambda$ is the chemical potential and $N$ is the particle number operator, while the angular brackets in Eq. (80) stand for the thermal average Abrikosov et al. (1975); Zagoskin (2014)
@@
$-\langle{\cal T}[V,{\hat{\Psi}}_{\mu\mu^{\prime}}](\tau){\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}(\tau^{\prime})\rangle,$ (85)
@@
^${\hat{\cal F}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=\langle{\cal T}[V,{\hat{\Psi}}_{\mu\mu^{\prime}}](\tau){\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}(\tau^{\prime})\rangle$ (87)
@@
^${\hat{\cal F}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})(\overleftarrow{\partial}_{\tau^{\prime}}$displaystyle-- ${\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}})$
@@
^${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=T\sum\limits_{n}e^{-\omega_{n}(\tau-\tau^{\prime})}{\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n}).$ (89)
@@
^$\hat{\cal R}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$ $=$ ^$\hat{\cal R}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$
@@
^$\hat{\cal R}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})=[\omega_{n}-{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}]^{-1}{\hat{\cal N}}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ (91)
@@
^${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$ $=$ ^${\hat{\cal R}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})+$
@@
with the kernel ${\hat{\cal K}}$ irreducible with respect to ^${\hat{\cal R}}^{0}$, or two-quasiparticle analog of the self-energy:
@@
Eqs. (LABEL:RTmatrixTLABEL:Tmatrix) set the complete and most general framework for the finite-temperature superfluid response in the 4$\times$4 block matrix form and in the domain of Matsubara frequencies, which is remarkably analogous to the non-superfluid and zero-temperature Dyson-BSEs. Further details depend essentially on the assumptions or direct knowledge about the ground and excited state structure, which can be expressed via expansions of quasiparticle operator strings. Some approximate solutions are discussed in the next subsections.
@@ II.2 Norm matrix, static kernel, and finite-temperature QRPA (FT-QRPA)
and operators $B,D$ and $G$ are listed in Appendix B. One can see, in particular, that the main diagonal of the norm matrix is formed by combinations of the $C$ and $C^{\dagger}$ operators, while the non-diagonal matrix elements are either zero (upper left block) or $A$-type operators (upper right and lower left blocks), or $C$ type ones but with one pair of permuted indices (lower right block). In the uncorrelated superfluid quasiparticle states, i.e., the FT-HFB states, all the non-diagonal contributions vanish, and the diagonal matrix elements are ${\cal N}^{[ii]}_{\mu\mu^{\prime}\nu\nu^{\prime}}=\delta_{\mu\mu^{\prime}\nu\nu^{\prime}}{\cal N}^{[ii]}_{\mu\mu^{\prime}}$,
@@
The minimally correlated superfluid interaction kernel is confined by the static contribution
${\hat{\cal K}}^{0{\cal N}}_{\mu\mu^{\prime}\nu\nu^{\prime}}=\frac{1}{4}\sum\limits_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal N}}^{-1}_{\mu\mu^{\prime}\gamma\gamma^{\prime}}{\hat{\cal T}}^{0}_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal N}}^{-1}_{\delta\delta^{\prime}\nu\nu^{\prime}},$ (108)
where ${\hat{\cal T}}^{0}$ without arguments stand for the matrix of the averaged double commutators
${\hat{\cal T}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}=-\langle[[V,{\hat{\Psi}}_{\mu\mu^{\prime}}],{\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}]\rangle,$ (109)
and constitute the FT-QRPA in the Dyson-BSE form:
^${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$ $=$ ^${\hat{\cal R}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})+$
$+$ $\frac{1}{4}\sum\limits_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal R}}^{0}_{\mu\mu^{\prime}\gamma\gamma^{\prime}}(\omega_{n}){\hat{\cal K}}^{0{\cal N}}_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal R}}_{\delta\delta^{\prime}\nu\nu^{\prime}}(\omega_{n}).$
The finite-temperature grand canonical ensemble averages for the FT-HFB matrix elements of Eq. (109) can be found using the commutators of Appendix B, and the results are given in Appendix B of Ref. Sommermann (1983).
@@ II.3 The dynamical kernel
The most general 4$\times$4 dynamical kernel, before including the norm factors,
${\cal K}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=\langle{\cal T}[V,\Psi^{[i]}_{\mu\mu^{\prime}}](\tau)[V,{\bar{\Psi}}^{[j]}_{\nu\nu^{\prime}}](\tau^{\prime})\rangle^{irr},$ (111)
@@
consists of products of the same set of commutators $[V,A]$, $[V,C]$ given in Appendix B and their Hermitian conjugates, and has in addition a $\tau$-dependence. We note that each term in Eq. (111) contains a product of eight quasiparticle operators, four at time $\tau$ and four at time $\tau^{\prime}$, i.e., correlated two-times four-quasiparticle propagator, contracted with two matrix elements of the residual interaction $H^{kl}$. As in the non-superfluid and zero-temperature cases, the appearance of higher-rank propagators generates a hierarchy of coupled EOMs for propagators of increasing rank. This means that to find viable solutions for realistic many-body systems, some reasonable approximations have to be applied to the dynamical kernel. Since the generic features of the EOM for the superfluid response, for instance, the interaction kernel decomposition and its operator ranks, are inherited by the finite-temperature theory from the zero-temperature one, we can apply a similar logic as was done, for instance, in Ref. Litvinova and Zhang (2022).
@@
The approximation confined by only one-fermion correlation functions is given by the complete factorization of the ${\cal K}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ into time-ordered quasiparticle operator pairs. This constitutes the finite-temperature superfluid analog of the second random phase approximation Yannouleas et al. (1985). Correlations that are more relevant to large fermionic systems with intermediate coupling should retain two-fermion correlation functions, which are known to form collective modes. Therefore, in the following, we will focus on such factorization of ${\cal K}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ into pairwise products of two-quasiparticle propagators. As in Ref. Litvinova and Zhang (2022) with zero-temperature formalism, mapping the two-quasiparticle correlated pairs to the superfluid phonons, introducing quasiparticle-vibration coupling and subsequently relaxing correlations in one propagator of each pair leads to the finite-temperature version of the leading-order superfluid NFT. Furthermore, the pairwise factorization enables a closed form of the Dyson-BSE with respect to the response function with the interesting possibility to iterate the dynamical kernel and, thus, include, in principle, arbitrarily complex 2n-quasiparticle configurations in the factorized form.
@@
The major difficulty at this step is the complexity of the residual interaction (14) and an even larger amount of components than in the case of the zero-temperature superfluid response. However, they have an analogous operator structure. The leading contributions to the dynamical kernel reside on the main diagonal of the $\hat{\cal K}^{r}$ matrix. The off-diagonal matrix elements represent various qPVC-induced ground state correlations of the resonant character. We will discuss the diagonal components in detail, and the off-diagonal ones can be found analogously. Let us consider the component ${\cal T}^{r[11]}(\tau-\tau^{\prime})$ and first concentrate on the terms associated with the $H^{31}$ matrix elements which played the leading role at zero temperature. As in the latter case, there are three types of operator products at $H^{31}$: (i) ’$AA$’-terms, containing operators $A$ and $A^{\dagger}$, (ii) ’$CC$’-terms with only $C$ and $C^{\dagger}$ operators and (iii) ’$AC$’-terms containing both operator types. The contribution of the first group of terms to ${\cal T}^{r[11]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ reads:
@@
×$\times\sum\limits_{\eta\eta^{\prime}\delta}\Bigl{[}\bigl{(}H^{31\ast}_{\eta\eta^{\prime}\nu\delta}(A^{\dagger}_{\nu^{\prime}\delta}A_{\eta\eta^{\prime}})(\tau^{\prime})$
@@
[…]
@@
×$\times\bigl{(}e^{-(\omega_{fi}+\omega_{f^{\prime}i^{\prime}})/T}-1\bigr{)},$ (114)
@@
where $\omega_{fi}=E_{f}-E_{i}$. Contracting the factorized CFs with the interaction matrix elements $H^{31}$, one finds that, as in the zero-temperature case, these contractions form two types which were classified as types (a) and (b) in Ref. Litvinova and Zhang (2022). A similar classification can be made at finite temperature, where only the amount of terms grows because of the additional contributions from the $\{{\cal U},{\cal V}\}$ amplitudes. In particular, in Eq. (113), the first terms in the last four lines are of type (b) and the second terms in those lines are of type (a). Collecting all the terms of type (a) to the dynamical kernel, the compact form of $\hat{\cal K}^{r[11]}$ in the energy domain can be recast as follows:
${\cal K}^{r[11]cc(a)}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})=\sum\limits_{\gamma\delta nm}p_{i_{n}}p_{i_{m}}\bigl{(}1-e^{-(\omega_{n}+\omega_{m})/T}\bigr{)}$ (115)
$\times$ $\Bigl{[}\frac{(\Gamma^{(11)n}_{\mu\gamma}{\cal X}^{m}_{\mu^{\prime}\gamma}+\Gamma^{(02){\bar{n}}\ast}_{\mu\gamma}{\cal U}^{m}_{\mu^{\prime}\gamma})({\cal X}^{m\ast}_{\nu^{\prime}\delta}\Gamma^{(11)n\ast}_{\nu\delta}+{\cal U}^{{\bar{m}}\ast}_{\nu^{\prime}\delta}\Gamma^{(02){\bar{n}}}_{\nu\delta})}{\omega_{k}-\omega_{n}-\omega_{m}}$
-- $\frac{(\Gamma^{(11)n\ast}_{\gamma\mu}{\cal Y}^{m\ast}_{\mu^{\prime}\gamma}+\Gamma^{(02){\bar{n}}}_{\gamma\mu}{\cal V}^{{\bar{m}}\ast}_{\mu^{\prime}\gamma})({\cal Y}^{m}_{\nu^{\prime}\delta}\Gamma^{(11)n}_{\delta\nu}+{\cal V}^{m}_{\nu^{\prime}\delta}\Gamma^{(02){\bar{n}}\ast}_{\delta\nu})}{\omega_{k}+\omega_{n}+\omega_{m}}\Bigr{]}$
-- $\cal{AS},$

### v4

@@ II.1.2 Finite temperature
${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=-\langle{\cal T}{\hat{\Psi}}_{\mu\mu^{\prime}}(\tau){\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}(\tau^{\prime})\rangle$ (80)
@@
[…]
@@
$-\langle{\cal T}[V,{\hat{\Psi}}_{\mu\mu^{\prime}}](\tau){\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}(\tau^{\prime})\rangle,\ \ \ \ \ \ \ \ \ \$(85)
@@
${\hat{\cal F}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=\langle{\cal T}[V,{\hat{\Psi}}_{\mu\mu^{\prime}}](\tau){\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}(\tau^{\prime})\rangle$ (87)
@@
${\hat{\cal F}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})(\overleftarrow{\partial}_{\tau^{\prime}}$ - ${\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}})$
@@
${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=T\sum\limits_{n}e^{-\omega_{n}(\tau-\tau^{\prime})}{\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n}).$ (89)
@@
$\hat{\cal R}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$ $=$ $\hat{\cal R}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$
@@
$\hat{\cal R}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})=[\omega_{n}-{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}]^{-1}{\hat{\cal N}}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ (91)
@@
${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$ $=$ ${\hat{\cal R}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})+$
@@
with the kernel $^{\hat{\cal K}}$ irreducible with respect to ${\hat{\cal R}}^{0}$, or two-quasiparticle analog of the self-energy:
@@
Eqs. (LABEL:RTmatrixTLABEL:Tmatrix) set the complete and most general framework for the finite-temperature superfluid response in the $\times4$ block matrix form and in the domain of Matsubara frequencies, which is remarkably analogous to the non-superfluid and zero-temperature Dyson-BSEs. Further details depend essentially on the assumptions or direct knowledge about the ground and excited state structure, which can be expressed via expansions of quasiparticle operator strings. Some approximate solutions are discussed in the next subsections.
@@ II.2 Norm matrix, static kernel, and finite-temperature QRPA (FT-QRPA)
and operators $B,D$ and $G$ are listed in Appendix B. One can see, in particular, that the main diagonal of the norm matrix is formed by combinations of the $C$ and $C^{\dagger}$ operators, while the non-diagonal matrix elements are either zero (upper left block) or AA-type operators (upper right and lower left blocks), or $C$ type ones but with one pair of permuted indices (lower right block). In the uncorrelated superfluid quasiparticle states, i.e., the FT-HFB states, all the non-diagonal contributions vanish, and the diagonal matrix elements are ${\cal N}^{[ii]\text{HFB}}_{\mu\mu^{\prime}\nu\nu^{\prime}}=\delta_{\mu\mu^{\prime}\nu\nu^{\prime}}{\cal N}^{[ii]}_{\mu\mu^{\prime}}$,
@@
The static contribution to the interaction kernel reads:
${\hat{\cal K}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}=\frac{1}{4}\sum\limits_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal N}}^{-1}_{\mu\mu^{\prime}\gamma\gamma^{\prime}}{\hat{\cal T}}^{0}_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal N}}^{-1}_{\delta\delta^{\prime}\nu\nu^{\prime}},$ (108)
where ${\hat{\cal T}}^{0}$ without arguments stands for the matrix of the averaged double commutators
${\hat{\cal T}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}=-\langle[[V,{\hat{\Psi}}_{\mu\mu^{\prime}}],{\hat{\bar{\Psi}}}_{\nu\nu^{\prime}}]\rangle.$ (109)
The minimally correlated superfluid interaction kernel is confined by the static contribution and, in combination with the HFB norm matrix, constitutes the FT-QRPA in the Dyson-BSE form:
${\hat{\cal R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})$ $=$ ${\hat{\cal R}}^{0;\text{HFB}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{n})+$
$+$ $\frac{1}{4}\sum\limits_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal R}}^{0;\text{HFB}}_{\mu\mu^{\prime}\gamma\gamma^{\prime}}(\omega_{n})\ {\hat{\cal K}}^{0;\text{QRPA}}_{\gamma\gamma^{\prime}\delta\delta^{\prime}}\ {\hat{\cal R}}_{\delta\delta^{\prime}\nu\nu^{\prime}}(\omega_{n}),$
where the index ”HFB” indicates the HFB approximation to the norm (107) in the uncorrelated propagators, and ”QRPA” marks the kernel which, in addition to the norm, uses the uncorrelated HFB variant of the generalized two-body density matrix defined in Appendix E and the well-known quasi-boson approximation Ring and Schuck (1980). The finite-temperature grand canonical ensemble averages for the FT-HFB matrix elements of Eq. (109) can be found using the commutators of Appendix B, and the results are given in Appendix B of Ref. Sommermann (1983). The correlated static kernel is presented in Appendix E in terms of its diagonal components, while the off-diagonal components can be found in a similar fashion.
@@ II.3 The dynamical kernel
The most general $\times4$ dynamical kernel is given by
${\tilde{\cal K}}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})=\langle{\cal T}[V,\Psi^{[i]}_{\mu\mu^{\prime}}](\tau)[V,{\bar{\Psi}}^{[j]}_{\nu\nu^{\prime}}](\tau^{\prime})\rangle^{irr},$ (111)
where the symbol $\tilde{\ }$” indicates that this is the kernel before including the norm factors, i.e.,
${\tilde{\cal K}}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ $=$ ${\cal T}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime}),$ (112)
${{\cal K}}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ $=$ $\frac{1}{4}\sum\limits_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{\cal N}}^{-1}_{\mu\mu^{\prime}\gamma\gamma^{\prime}}{\tilde{\cal K}}^{r[ij]}_{\gamma\gamma^{\prime}\delta\delta^{\prime}}(\tau-\tau^{\prime}){\hat{\cal N}}^{-1}_{\delta\delta^{\prime}\nu\nu^{\prime}}.$
@@
It consists of products of the same set of commutators $[V,A]$, $[V,C]$ given in Appendix B and their Hermitian conjugates, and has in addition a $\tau$-dependence. We note that each term in Eq. (111) contains a product of eight quasiparticle operators, four at time $\tau$ and four at time $\tau^{\prime}$, i.e., correlated two-times four-quasiparticle propagator, contracted with two matrix elements of the residual interaction $H^{kl}$. As in the non-superfluid and zero-temperature cases, the appearance of higher-rank propagators generates a hierarchy of coupled EOMs for propagators of increasing rank. This means that to find viable solutions for realistic many-body systems, some reasonable approximations have to be applied to the dynamical kernel. Since the generic features of the EOM for the superfluid response, for instance, the interaction kernel decomposition and its operator ranks, are inherited by the finite-temperature theory from the zero-temperature one, we can apply a similar logic as was done, for instance, in Ref. Litvinova and Zhang (2022).
@@
The approximation confined by only one-fermion correlation functions is given by the complete factorization of the ${\cal K}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ into time-ordered quasiparticle operator pairs. This constitutes the finite-temperature superfluid analog of the second random phase approximation Yannouleas et al. (1985). Correlations that are more relevant to large fermionic systems with intermediate coupling should retain two-fermion correlation functions, which are known to form collective modes. Therefore, in the following, we will focus on such factorization of ${\cal K}^{r[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ into pairwise products of two-quasiparticle propagators. As in Ref. Litvinova and Zhang (2022) with zero-temperature formalism, mapping the two-quasiparticle correlated pairs to the superfluid phonons, introducing quasiparticle-vibration coupling and subsequently relaxing correlations in one propagator of each pair leads to the finite-temperature version of the leading-order superfluid NFT. Furthermore, the pairwise factorization enables a closed form of the Dyson-BSE with respect to the response function, with the interesting possibility to iterate the dynamical kernel and, thus, include, in principle, arbitrarily complex 2n-quasiparticle configurations in the factorized form.
@@
The major difficulty at this step is the complexity of the residual interaction (14) and an even larger number of components than in the case of the zero-temperature superfluid response. However, they have an analogous operator structure. The leading contributions to the dynamical kernel reside on the main diagonal of the $\hat{\cal K}^{r}$ matrix. The off-diagonal matrix elements represent various qPVC-induced ground state correlations of the resonant character. We will discuss the diagonal components in detail, and the off-diagonal ones can be found analogously. Let us consider the component ${\cal T}^{r[11]}(\tau-\tau^{\prime})$ and first concentrate on the terms associated with the $H^{31}$ matrix elements, which played the leading role at zero temperature. As in the latter case, there are three types of operator products at $H^{31}$: (i) ’AAAA’-terms, containing operators AA and $A^{\dagger}$, (ii) ’CCCC’-terms with only $C$ and $C^{\dagger}$ operators and (iii) ’ACAC’-terms containing both operator types. The contribution of the first group of terms to ${\cal T}^{r[11]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau-\tau^{\prime})$ reads:
@@
$\times\sum\limits_{\eta\eta^{\prime}\delta}\Bigl[\bigl(H^{31\ast}_{\eta\eta^{\prime}\nu\delta}(A^{\dagger}_{\nu^{\prime}\delta}A_{\eta\eta^{\prime}})(\tau^{\prime})$
@@
$[{\cal R}^{[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}{\cal R}^{[kl]}_{\eta\eta^{\prime}\rho\rho^{\prime}}](\omega_{k})=\int_{-1/T}^{1/T}d\tau e^{\omega_{k}\tau}{\cal R}^{[ij]}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\tau){\cal R}^{[kl]}_{\eta\eta^{\prime}\rho\rho^{\prime}}(\tau)$
@@
$\times\bigl(e^{-(\omega_{fi}+\omega_{f^{\prime}i^{\prime}})/T}-1\bigr),\ \ \ \ \ \ \ \ \ \ \ \ \ \$(116)
@@
where $\omega_{fi}=E_{f}-E_{i}$. Contracting the factorized CFs with the interaction matrix elements $H^{31}$, $H^{40}$, and $H^{22}$, one finds that, as in the zero-temperature case, these contractions form two types which were classified as types (a) and (b) in Ref. Litvinova and Zhang (2022). A similar classification can be made at finite temperature, where only the number of terms grows because of the additional contributions from the $\{{\cal U},{\cal V}\}$ amplitudes. In particular, in Eq. (115), the first terms in the last four lines are of type (b) and the second terms in those lines are of type (a). Collecting all the terms of type (a) in the dynamical kernel and dropping the reducible terms belonging to the constant and one-body sectors of the Hamiltonian, the compact form of $\tilde{\cal K}^{r[11]}$ and $\tilde{\cal K}^{r[33]}$ (again, before the inclusion of the norm factors) in the energy domain can be recast as follows:
${\tilde{\cal K}}^{r[11]cc(a)}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})=\sum_{\gamma\delta nm}p_{i_{n}}p_{i_{m}}(1-e^{-(\omega_{n}+\omega_{m})/T})$
$\times\Bigl[\frac{(\Gamma^{(11)n}_{\mu\gamma}{\cal X}^{m}_{\mu^{\prime}\gamma}+\Gamma^{(20)n}_{\mu\gamma}{\cal V}^{m}_{\mu^{\prime}\gamma})({\cal X}^{m\ast}_{\nu^{\prime}\delta}\Gamma^{(11)n\ast}_{\nu\delta}+{\cal V}^{m\ast}_{\nu^{\prime}\delta}\Gamma^{(20)n\ast}_{\nu\delta})}{\omega_{k}-\omega_{n}-\omega_{m}}$
$-\frac{(\Gamma^{(11){\bar{n}}}_{\mu\gamma}{\cal X}^{\bar{m}}_{\mu^{\prime}\gamma}+\Gamma^{(20){\bar{n}}}_{\mu\gamma}{\cal V}^{\bar{m}}_{\mu^{\prime}\gamma})({\cal X}^{{\bar{m}}\ast}_{\nu^{\prime}\delta}\Gamma^{(11){\bar{n}}\ast}_{\nu\delta}+{\cal V}^{{\bar{m}}\ast}_{\nu^{\prime}\delta}\Gamma^{(20){\bar{n}}\ast}_{\nu\delta})}{\omega_{k}+\omega_{n}+\omega_{m}}\Bigr]$
$-\cal{AS},\ \ \ \ \ \ \$(117)
${\tilde{\cal K}}^{r[33]cc(a)}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})=\sum_{\gamma\delta nm}p_{i_{n}}p_{i_{m}}(1-e^{-(\omega_{n}+\omega_{m})/T})$
$\times\Bigl[\frac{(\Gamma^{(11)n}_{\gamma\mu}{\cal U}^{m}_{\gamma\mu^{\prime}}+\Gamma^{(20)n}_{\mu^{\prime}\gamma}{\cal Y}^{m}_{\mu\gamma})({\cal U}^{m\ast}_{\delta\nu^{\prime}}\Gamma^{(11)n\ast}_{\delta\nu}+{\cal Y}^{m\ast}_{\nu\delta}\Gamma^{(20)n\ast}_{\nu^{\prime}\delta})}{\omega_{k}-\omega_{n}-\omega_{m}}$
$-\frac{(\Gamma^{(11){\bar{n}}}_{\gamma\mu}{\cal U}^{\bar{m}}_{\gamma\mu^{\prime}}+\Gamma^{(20){\bar{n}}}_{\mu^{\prime}\gamma}{\cal Y}^{\bar{m}}_{\mu\gamma})({\cal U}^{{\bar{m}}\ast}_{\delta\nu^{\prime}}\Gamma^{(11){\bar{n}}\ast}_{\delta\nu}+{\cal Y}^{{\bar{m}}\ast}_{\nu\delta}\Gamma^{(20){\bar{n}}\ast}_{\nu^{\prime}\delta})}{\omega_{k}+\omega_{n}+\omega_{m}}\Bigr]$
$-\cal{AS}^{\dagger},\ \ \ \ \ \$(118)
$=\sum_{\gamma\delta nm}p_{i_{n}}p_{i_{m}}(1-e^{-(\omega_{n}+\omega_{m})/T})$
$\times\Bigl[\frac{(\Gamma^{(11){\bar{n}}\ast}_{\mu\gamma}{\cal U}^{{\bar{m}}\ast}_{\mu^{\prime}\gamma}-\Gamma^{(20){\bar{n}}\ast}_{\mu\gamma}{\cal Y}^{{\bar{m}}\ast}_{\mu^{\prime}\gamma})({\cal U}^{{\bar{m}}}_{\nu^{\prime}\delta}\Gamma^{(11){\bar{n}}}_{\nu\delta}-{\cal Y}^{{\bar{m}}}_{\nu^{\prime}\delta}\Gamma^{(20){\bar{n}}}_{\nu\delta})}{\omega_{k}-\omega_{n}-\omega_{m}}$
$-\frac{(\Gamma^{(11){n}\ast}_{\mu\gamma}{\cal U}^{{m}\ast}_{\mu^{\prime}\gamma}-\Gamma^{(20){n}\ast}_{\mu\gamma}{\cal Y}^{{m}\ast}_{\mu^{\prime}\gamma})({\cal U}^{{m}}_{\nu^{\prime}\delta}\Gamma^{(11){n}}_{\nu\delta}-{\cal Y}^{{m}}_{\nu^{\prime}\delta}\Gamma^{(20){n}}_{\nu\delta})}{\omega_{k}+\omega_{n}+\omega_{m}}\Bigr]$
$-\cal{AS}^{\dagger},\ \ \ \ \ \$(119)